# Capturing strong correlation effects on a quantum annealer: calculation of avoided crossing in the H4 molecule using the quantum annealer eigensolver

### v1

@@ Abstract
We broaden the scope of the Quantum Annealer Eigensolver (QAE) algorithm, an underexplored noisy intermediate scale quantum (NISQ) era approach for calculating atomic and molecular properties, to predict avoided crossings, where strong correlation effects are at play. For this purpose, we consider the classic example of the H4 molecule in a rectangular geometry. Our results are obtained on the 5000-qubit D-Wave Advantage system 4.1 quantum computer. We benchmark our quantum annealing results with full configuration interaction (FCI) as well as with those obtained using simulated annealing. We find that we can predict avoided crossings within about 1.1% of the FCI value on real quantum hardware. We carry out analyses on the effect of the number of shots, anneal time, and the choice of Lagrange multiplier on our obtained results. Since the QAE algorithm provides information on the wave function as its output, we also check the quality of the computed wave function by calculating the fidelity, and find it to be 99.886%. Finally, we qualitatively discuss the strengths and weaknesses of the QAE algorithm relative to its gate-based NISQ algorithm counterpart, the celebrated Variational Quantum Eigensolver. Our work contributes to the existing body of literature on QAE by demonstrating that high-quality results can be achieved on noisy hardware.
@@ I Introduction
Due to its general purpose nature, Quantum Annealer Eigensolver (QAE), which is a noisy intermediate scale quantum (NISQ) era quantum annealing algorithm [1] built on the variational principle, has been widely applied to a variety of problems. They include molecular vibrational spectra [1], complex eigenvalue problems [2], energy calculations for molecular electronic states [3, 4], relativistic calculations of fine structure splitting in highly charged atomic ions [5], simulating particle physics [6], and lattice gauge theories [7]. In one of these works, the authors demonstrated that QAE could capture the fine structure splittings in highly charged atomic ions with unprecedented accuracy on quantum annealer hardware [5]. This aspect is especially noteworthy, since in gate-based NISQ hardware, noise poses significant problems, and thus obtaining results of comparable quality using similar approaches such as the variational quantum eigensolver (VQE) is extremely challenging. However, despite its general scope and better resilience to noise, the QAE algorithm has been vastly underexplored relative to VQE, which has received a lot of attention both on quantum hardware (for example, see Refs. [8, 9, 10, 11, 12, 13, 14] and simulation fronts (Refs. [15, 16, 17, 18, 19, 20, 21, 22]).
The versatility and robustness of QAE to quantum hardware errors as well as dearth in the range of applications in the context of quantum chemistry naturally prompts further exploration of obtaining other molecular properties using the algorithm. In particular, it is interesting and timely to study the performance of QAE in domains where predicting strong correlation effects, one of the cornerstones of quantum many-body theoretic calculations applied to quantum chemistry, are at play. This would be a first step towards mirroring the tremendous efforts that have been dedicated towards capturing strong correlation effects in the VQE framework [23, 24, 25, 26, 27, 28].
@@
Figure 1: Molecular orbital diagram for H4 across its potential energy curve, for the cases (a) $\alpha<a$, (b) $\alpha\approx a$, and (c) $\alpha>a$, where $\alpha$ denotes the distance between two $H_{2}$ sub-systems that constitute H4, whereas $a$ is the bond length of the $H_{2}$ molecule. In the sub-figures, $a_{g}$, $b_{1u}$, etc refer to the irreducible representations of the $D_{2h}$ point group, while $\Delta\varepsilon_{i}$ is the difference in energy between the HOMO and LUMO across the PEC. HOMO and LUMO are the highest occupied and the lowest unoccupied molecular orbitals respectively.
In this work, we extend the range of applicability of the QAE algorithm into the strong correlation regime by carrying out avoided crossings (AC) calculations on the well-known prototypical system: H4 in a rectangular geometry [29]. We expand the many-body wavefunction in a basis of configuration state functions (CSFs) and recast the energy functional as an Ising Hamiltonian. We then determine the ground state energy and the excited state energy of interest to us using three approaches: quantum annealing (for which we employ the QAE algorithm; on the D-Wave Advantage machine), simulated annealing (for which we use simulated annealing eigenvoler (SAE) approach), and the graphical unitary group approach full configuration interaction (GUGA-FCI; which we shall hereafter shorten to FCI for brevity) provided by GAMESS-US [30] for benchmarking our results. We also carry out an analysis of the ‘knobs’ of the QAE algorithm, such as the effect of the number of shots, the anneal time, and the choice of the Lagrange parameter on our AC results. Finally, since QAE and VQE are both NISQ algorithms, we compare the two approaches.
The rest of the sections are organized as follows: Section II presents the relevant theory and methodology, while Section III discusses our findings. In Section IV we compare QAE with VQE and conclude with Section V.
@@ II.1 Quantum annealing
QA is a meta-heuristic that leverages quantum fluctuations, and is designed to solve computationally hard optimization problems, typically by encoding them into the ground state of a Hamiltonian  [31, 32, 33]. The underlying principle behind QA is the adiabatic theorem. Originally proposed by Born and Fock in 1928 [34], the adiabatic theorem can be encapsulated as follows: Given a time-dependent Hamiltonian, $H(t)=\bigg{(}1-\frac{t}{T}\bigg{)}H_{I}+\frac{t}{T}H_{F};\ t\in[0,T]$, which is initially $H_{I}$ at $t=0$ and subsequently $H_{F}$ at some later time, $t=T$, and given that the system starts in the ground state of $H_{I}$, the state of the system is guaranteed to remain in the ground state throughout its time evolution as long as the Hamiltonian is varied sufficiently slowly. Thus, the state of the system is in the ground state of $H_{F}$ at $t=T$. QA involves preparing the ground state of an easily constructible Hamiltonian, $H_{I}$, which is typically chosen to be the transverse field Hamiltonian, and then gradually transforming it into another Hamiltonian, $H_{F}$, generally chosen to be the Ising Hamiltonian whose ground state encodes the solution to the problem. As one gradually goes from $H_{I}$ to $H_{F}$, the system transitions smoothly from the ground state of $H_{I}$ to the ground state of $H_{F}$. Since QA is naturally suited to finding global minima of energy landscapes, we use it to obtain the ground state energies of electronic structure Hamiltonians, which are expressed in the Ising form. We use the QAE algorithm for this purpose [1].
@@ II.2 The QAE algorithm: a primer
The QAE algorithm solves the eigenvalue equation $H\ket{\Psi}=E\ket{\Psi}$ by transforming it into an energy minimization problem. The energy functional considered is the expectation value of the Hamiltonian $H$ with respect to the unknown state |$\ket{\Psi}$. To avoid trivial solutions, the normalisation constraint $\langle\Psi|\Psi\rangle-1=0$ is enforced by the inclusion of a Lagrange multiplier $\lambda$ into the energy functional. The modified formulation after dropping the irrelevant constant is given by
@@
Assuming an ansatz of the form |$\ket{\Psi}=\sum^{N}_{i=1}c_{i}\ket{\Phi_{i}}$ where {|$\{\ket{\Phi_{i}}\}$ represent a set of $N$ known basis functions, the QAE algorithm aims to determine the unknown coefficients $\{c_{i}\}$ with $c_{i}\in[-1,1]$. These coefficients are encoded into $K$ bits using the fixed point encoding scheme. This approach converts the problem from a continuous optimization over the variables $\{c_{i}\}$
$\epsilon(\vec{c},\lambda)$ $=\sum_{i,j=1}^{N}c_{i}c_{j}H_{ij}-\lambda\sum_{i=1}^{N}c_{i}^{2}.$ (2)
into a discrete optimization problem over the binary variables $\{q_{i}\}$ where $q_{i}\in\{-1,1\}$ making it suitable for implementation on modern day quantum annealers. In the above equation, $H_{ij}$ refers to the matrix elements of the Hamiltonian.
The pseudocode below summarizes the QAE implementation employed.
@@ II.4 Computational details
To generate the molecular Hamiltonian, we use the GUGA-FCI approach available in the GAMESS-US program [30]. There are 8 CSFs for our 4 electron 4 orbital active space. This corresponds to an $80\times 80$ all-to-all connected QUBO problem with $K=10$.
@@ III Results and Discussions
We now discuss our results. Figure 2 presents our main result (with accompanying data provided in Table 1), where we compare the predicted avoided crossing from QAE with SAE and FCI, with 1000 shots and 10 repetitions. From the figure, we see that QAE, SAE, and FCI yield an energy difference (at the AC geometry, that is, 2 Bohr) of 0.17195, 0.15037, and 0.15169 Hartree (Ha) respectively. Thus, QAE is able to predict AC in the chosen system to within about 13 percent of the FCI value. The plot also shows that SAE is clearly in better agreement with the FCI values than QAE with both the ground and excited state energies at all of the geometries considered. Furthermore, the quality of the predicted excited state energy is relatively poor for the AC geometry using QAE (as we shall discuss later, the quality of results from QAE is still substantially better than those from a VQE computation with the same parameters). We can backtrack the performance of QAE to the several input parameters that go into obtaining a QAE result, such as (but not limited to) the number of shots and number of repetitions, anneal time, and choice of $\lambda$ range. We devote the subsequent paragraphs to analyzing the effect of these three parameters on the AC results.
@@
Figure 4: Comparison of CI coefficients for the ground state obtained using QAE with exact values of the coefficients from FCI (denoted as a brown triangle) for 1000 and 5000 shots, both with 30 repetitions, at the AC geometry. Each of the pink (1000 shots) and green (5000 shots) circles denotes the result from a repetition.
@@
We begin by carrying out AC calculations with 1000 shots and 30 repetitions. As Figure 2 shows, increasing the number of repetitions drastically improves the quality of the excited state energy (the difference with respect to the FCI value) improves from 25.8 mHa to 7 mHa and improves the agreement of the ground state energy with the FCI value to 4.4 mHa (from 5.6 mHa). The AC value improves from 20.2 mHa for 10 repetitions to 2.6 mHa for 30 repetitions. We extend the analysis further by increasing the number of repetitions all the way to 100, but find that the AC value remains unchanged. Thus, we instead increase the number of shots to 5000 while keeping the number of repetitions fixed at 30. Figure 3 presents our findings, with the accompanying data given in Table 1. The results indicate that as expected, one obtains better agreement with the FCI value for the AC with larger number of shots and repetitions. Furthermore, we see that the QAE result now is in much better agreement with the FCI value, differing only by 1.14 percent (5000 shots, 30 repetitions), as opposed to the 1000 shots 10 repetitions value of 13.35 percent and the 1000 shots 30 repetitions value of 3.2 percent. These percentages correspond to an energy difference ($E_{AC,FCI}-E_{AC,QAE}$) of 1.72 mHa, 20.2 mHa, and 2.6 mHa respectively. The values of AC obtained with 5000 shots and 30 repetitions, 1000 shots and 10 repetitions, and 1000 shots and 30 repetitions are 0.15342, 0.17195, and 0.15654 Ha respectively.
We turn our attention to Figure 4, where we plot the CI coefficients that we obtained from the ground state QAE calculation versus the coefficient index. We benchmark the computed CI coefficients employing QAE with those obtained from FCI. We carry out the exercise for both the 1000 and the 5000 shot cases, each with 30 repetitions. The purpose of the plot is to check the quality of the coefficients, which in turn indicates the quality of the predicted wave function. We see that the coefficients are all in reasonable agreement with their FCI counterparts. While the plot itself seems to yield almost similar set of coefficients by using 5000 shots over 1000 shots, this seemingly insignificant difference between their results is reflected in the subsequent ground state energy calculations as seen in the preceding paragraphs. Unsurprisingly, we also see the chemistry at play, where two coefficients are almost equally important (unlike the single reference scenario where only the HF state dominates). We also check the fidelity as well as the difference between AC predicted using QAE and FCI, and find them to be 99.735% and 2.6mHa respectively for QAE calculation with 1000 shots and 30 repetitions, whereas they improve to 99.886% and 1.72 mHa respectively when we employ 5000 shots and 30 repetitions. Figure 5 presents scatter plots between fidelity ($\mathcal{F}$) and $\Delta E$, the difference between the QAE and the FCI ground state energy results. The fidelity is given by $|\langle\Psi_{QAE}|\Psi_{FCI}\rangle|^{2}$, where $|\Psi_{QAE}\rangle$ is the wave function constructed using the CI coefficients that QAE outputs, while $|\Psi_{FCI}\rangle$ refers to the FCI wave function. We carry out linear regression analysis on our two data sets, and find that the Pearson coefficient, $R$, is weakly correlated (0.115) for the 1000 shots case but is weakly anti-correlated (-0.22) in the 5000 shots scenario. Ideally, we would expect that as we supply more shots, the strength of anti-correlation between the two quantities, $\mathcal{F}$ and $\Delta E$, increases, based on the inequality $\mathcal{F}\geq 1-\frac{\Delta E}{E_{AC,true}}$, where $E_{AC,true}$ is the FCI value for the energy at the AC. Our data supports this expectation. Furthermore, we find the best-case fidelities and ground state energies that we obtain using QAE for both the 1000 and the 5000 shots cases (99.735% and 99.886% respectively) obey the aforementioned inequality relation.
@@
Figure 5: Scatter plot involving fidelity and $\Delta E$ values. The former quantity refers to $|\langle\Psi_{QAE}|\Psi_{FCI}\rangle|^{2}$, where $|\Psi_{QAE}\rangle$ is the wave function constructed out of the CI coefficients obtained using QAE and $|\Psi_{FCI}\rangle$ is the FCI wave function. On the other hand, the latter is the difference between the QAE and the FCI ground state energy results.
@@
Finally, we focus on the choice of $\lambda$ range for the problem. As mentioned in an earlier paragraph in Section II, we picked the $\lambda$ range that gave the best SAE result. For the purposes of this work, we pick a range of 0.1 Ha. However, since SAE and QAE are different approaches to search for a desired solution, we pick three such ranges of 0.1 Ha each: a range containing the HF value (-1.77 Ha), which is -1.8 to -1.7 Ha, one that is immediately below and another that is immediately above it. We term the three ranges as A, B, and C respectively. We note that the range we picked for our main calculations was the one immediately above the HF value, that is, range C. Figure 7 presents our results justifying this choice. We observe that our choice of $\lambda$ range from SAE calculations were sufficiently good within a calculation involving 30 repetitions.

### v4

# Quantum annealing eigensolver as a NISQ era tool for probing strong correlation effects in quantum chemistry

@@ Abstract
The quantum-classical hybrid variational quantum eigensolver (VQE) algorithm is arguably the most popular noisy intermediate-scale quantum (NISQ) era approach to quantum chemistry. We consider the underexplored quantum annealing eigensolver (QAE) algorithm as a worthy alternative. We use a combination of numerical calculations for a system where strong correlation effects dominate, and conclusions drawn from our preliminary scaling analysis for QAE and VQE to make the case for QAE as a NISQ era contender to VQE for quantum chemistry. For the former, we pick the representative example of computing avoided crossings in the H4 molecule in a rectangular geometry, and demonstrate that we obtain results to within about 1.2% of the full configuration interaction value on the D-Wave Advantage system 4.1 hardware. We carry out analyses on the effect of the number of shots, anneal time, and the choice of Lagrange multiplier on our obtained results. Following our numerical results, we carry out a detailed yet preliminary analysis of the scaling behaviours of both the QAE and the VQE algorithms to assess the competency of the former for NISQ era chemistry.
@@ I Introduction
The Variational Quantum Eigensolver (VQE) algorithm is the most well-known and arguably the most widely used NISQ era approach to quantum chemistry. Several works in literature have employed the algorithm, both on quantum hardware (for example, see Refs. [1, 2, 3, 4, 5, 6, 7] and simulation fronts (Refs. [8, 9, 10, 11, 12, 13, 14, 15]). However, in gate-based NISQ hardware, noise poses significant problems, and thus obtaining results of high quality using VQE (or for that matter, any other gate-based algorithm applied to chemistry) is extremely challenging. On the other hand, on the quantum annealing front, the Quantum Annealer Eigensolver (QAE) [16], which is also a NISQ era quantum-classical hybrid approach and which is also built on the variational principle, has proven to yield highly accurate results [17]. Furthermore, due to its general purpose nature, QAE has been widely applied to a variety of problems. They include molecular vibrational spectra [16], complex eigenvalue problems [18], energy calculations for molecular electronic states [19, 20], relativistic calculations of fine structure splitting in highly charged atomic ions [17], simulating particle physics [21], and lattice gauge theories [22]. However, despite its general scope and better resilience to noise, the QAE algorithm has been vastly underexplored relative to VQE.
The versatility and robustness of QAE to quantum hardware errors as well as dearth in the range of applications in the context of quantum chemistry naturally prompts further exploration of obtaining molecular properties using the algorithm. In particular, it is interesting and timely to study the performance of QAE in domains where predicting strong correlation effects, one of the cornerstones of quantum many-body theoretic calculations applied to quantum chemistry, are at play. This would be a step towards mirroring the tremendous efforts that have been dedicated towards capturing strong correlation effects in the VQE framework [23, 24, 25, 26, 27, 28].
While other quantum annealing algorithms for quantum chemistry do exist, such as the Xia–Bian–Kais (XBK) transformation  [29, 30] and the Qubit Coupled Cluster (QCC) [31, 32] method, they have practical limitations. For instance, the XBK approach incurs a significant qubit overhead due to the need to encode a k-local electronic structure Hamiltonian (initially represented as a weighted sum of tensor product of all Pauli operators) into a form involving only Pauli $Z$ operators. The qubit count increases further due to quadratization required to transform it into a 2-local Ising Hamiltonian. In contrast, QAE minimises an energy functional expressed as a double sum over Hamiltonian matrix elements. Post encoding, this functional naturally results in a 2-local form compatible with the quantum annealer (see Eq. (4) below) thus avoiding the need for quadratization, the tradeoff here being that the encoded problem always exhibits a full connectivity among qubits. The QCC method on the other hand, although requiring fewer qubits than XBK [33], relies heavily on a classical optimizer to minimise the energy functional which can involve parameters scaling exponentially with the number of entanglers used to capture electron correlation [32]. Moreover, it was highlighted in [32] that the QCC method tends to exhibit the issue of convergence to different local minima. This problem is particularly severe for strongly correlated systems.
@@
Figure 1: Molecular orbital diagram for H4 across its potential energy curve, for the cases (a) $\alpha<a$, (b) $\alpha\approx a$, and (c) $\alpha>a$, where $\alpha$ denotes the distance between two H2 sub-systems that constitute H4, whereas $a$ is the bond length of the H2 molecule. In the sub-figures, $a_{g}$, $b_{1u}$, etc refer to the irreducible representations of the $D_{2h}$ point group, while $\Delta\varepsilon_{i}$ is the difference in energy between the HOMO and LUMO across the PEC. HOMO and LUMO are the highest occupied and the lowest unoccupied molecular orbitals respectively.
In this work, we begin by carrying out avoided crossings (AC) calculations on the well-known prototypical system: H4 in a rectangular geometry [34], to showcase the strength of QAE in predicting physical effects in the strong correlation regime. We expand the many-body wave function in a basis of configuration state functions (CSFs) and recast the energy functional as an Ising Hamiltonian. We then determine the ground state energy and the excited state energy of interest to us using three approaches: quantum annealing (for which we employ the QAE algorithm; on the D-Wave Advantage machine), simulated annealing (for which we use simulated annealing eigenvoler (SAE) approach), and the graphical unitary group approach full configuration interaction (GUGA-FCI; which we shall hereafter shorten to FCI for brevity) provided by GAMESS-US [35] for benchmarking our results. We also carry out an analysis of the ‘knobs’ of the QAE algorithm, such as the effect of the number of shots, the anneal time, and the choice of the Lagrange parameter on our AC results. For completeness, we also carry out VQE calculation for ground state energy estimation on the IBM superconducting devices.
Our numerical computations are followed by a preliminary analysis on the scaling behaviours of both the QAE and the VQE approaches (classical as well as quantum resources), with the goal of comparing the two. To that end, we consider the scaling of each of the steps of both the algorithms under some assumptions to arrive at the final scaling estimates.
The rest of the sections are organized as follows: Section II presents the relevant theory and methodology, while Section III discusses our findings from numerical calculations. In Section IV, we study the scaling behaviours of QAE and VQE, and conclude with Section V.
@@ II.1 Quantum annealing
Quantum annealing (QA) is a meta-heuristic that leverages quantum fluctuations, and is designed to solve computationally hard optimization problems, typically by encoding them into the ground state of a Hamiltonian  [36, 37, 38]. QA involves preparing the ground state of an easily constructible Hamiltonian, $H_{I}$, which is typically chosen to be the transverse field Hamiltonian, and then gradually transforming it into another Hamiltonian, $H_{F}$, generally chosen to be the Ising Hamiltonian whose ground state encodes the solution to the problem. As one gradually goes from $H_{I}$ to $H_{F}$ according to $H(t)=\bigg{(}1-\frac{t}{T}\bigg{)}H_{I}+\frac{t}{T}H_{F};\ t\in[0,T]$, the system transitions smoothly from the ground state of $H_{I}$ to the ground state of $H_{F}$. We add that that while the adiabatic model described above is the theoretical model underlying QA, the latter, in practice, operates far from the adiabatic limit. In this work we use QAE  [16] to map the electronic structure Hamiltonian, typically a k-local operator expressed as a weighted sum of Pauli strings, into an equivalent Ising form. We subsequently use QA to find its ground state.
@@ II.2 The QAE algorithm: a primer
The QAE algorithm solves the eigenvalue equation $H\ket{\Psi}=E\ket{\Psi}$ by transforming it into an energy minimization problem. The energy functional considered is the expectation value of the Hamiltonian $H$ with respect to an unknown state |$\ket{\Psi}$. To avoid trivial solutions, the normalisation constraint $\langle\Psi|\Psi\rangle-1=0$ is enforced by the inclusion of a Lagrange multiplier $\lambda$ into the energy functional. The modified formulation after dropping the irrelevant constant is given by
@@
Assuming an ansatz of the form |$\ket{\Psi}=\sum^{d_{CI}}_{i=1}c_{i}\ket{\Phi_{i}}$ where {|$\{\ket{\Phi_{i}}\}$ represent a set of $d_{CI}$ known basis functions, the QAE algorithm aims to determine the unknown coefficients $\{c_{i}\}$ with $c_{i}\in[-1,1]$ in the energy functional
$\epsilon(\vec{c},\lambda)=\sum_{i,j=1}^{d_{CI}}c_{i}c_{j}H_{ij}-\lambda\sum_{i=1}^{d_{CI}}c_{i}^{2}.$ (2)
We subsequently use the fixed point encoding scheme where each coefficient $c_{i}$ is encoded using K qubits $q^{i}_{K}\in\{-1,1\}$ in the following manner
$c_{i}=\sum^{K-1}_{\alpha=1}2^{\alpha-K}q^{i}_{\alpha}-q^{i}_{K}$ (3)
This approach converts the problem from a continuous optimization over the variables $\{c_{i}\}$ to a discrete optimization problem over the binary variables $\{q_{i}\}$ i.e.
$\epsilon(\vec{q},\lambda)$ $=\sum_{i,j=1}^{d_{CI}}\Big{(}\sum_{\alpha,\beta=1}^{K-1}2^{\alpha+\beta-2K}q^{i}_{\alpha}q^{j}_{\beta}+$ (4)
$2^{\alpha+1-K}q^{i}_{\alpha}\pm 1\Big{)}\Big{(}H_{ij}-\lambda\delta_{ij}\Big{)}$
making it suitable for implementation on current quantum annealers. In the above equation, $H_{ij}\equiv\langle\Phi_{i}|H|\Phi_{j}\rangle$ and $\delta_{ij}$ refers to the Kronecker delta function. The $\pm$ sign arises from the product $q^{i}_{K}q^{j}_{K}$ which can be 1 or $-1$. The pseudocode (see Algorithm II.2) summarizes the QAE implementation employed.
@@ II.4 Computational details
To generate the CI matrix elements, $H_{ij}$, we use the GUGA-FCI approach available in the GAMESS-US program [35]. There are 8 CSFs for our 4 electron 4 orbital active space. This corresponds to an $80\times 80$ all-to-all connected QUBO problem with $K=10$.
@@ III Results and Discussions
III.1 QAE results for avoided crossings
We now discuss our results. Figure 2 presents our results from annealing (with accompanying data provided in Table 1), where we compare the predicted avoided crossing from QAE with SAE and FCI, with 1000 shots and 10 repetitions. From the figure, we see that QAE, SAE, and FCI yield an energy difference (at the AC geometry, that is, 2 Bohr) of 0.17195, 0.15037, and 0.15169 Hartree (Ha) respectively. Thus, QAE is able to predict AC in the chosen system to within about 13 percent of the FCI value. The plot also shows that SAE is clearly in better agreement with the FCI values than QAE with both the ground and excited state energies at all of the geometries considered. A major bottleneck in QAE which is otherwise absent in SAE is the requirement for embedding. The QAE formulation results in a problem graph which is inherently fully connected regardless of the specific problem instance. Embedding thus refers to the process of mapping this fully connected graph onto sparsely connected qubits of the quantum annealer (here Pegasus topology [45] of D-Wave). This process incurs a significant qubit overhead and thus more error due to the increase of the search space (given by the physical Hilbert space dimension) while the logical solution space remains constant. SAE on the other hand is less error prone since classical computers are more robust to errors. Furthermore, the quality of the predicted excited state energy is relatively poor for the AC geometry using QAE (as we shall discuss later, the quality of results from QAE is still substantially better than those from a VQE computation with the same parameters). We can backtrack the performance of QAE to the several input parameters that go into obtaining a QAE result, such as (but not limited to) the number of shots and number of repetitions, anneal time, and choice of $\lambda$ range. We devote the subsequent paragraphs to analyzing the effect of these three parameters on the AC results.
@@ III.1 QAE results for avoided crossings
Figure 4: Comparison of configuration interaction coefficients for the ground state obtained using QAE with exact values of the coefficients from FCI (denoted as a brown triangle) for 1000 and 5000 shots, both with 30 repetitions, at the AC geometry. Each of the pink (1000 shots) and green (5000 shots) circles denotes the result from a repetition.
@@
We begin by carrying out AC calculations with 1000 shots and 30 repetitions. As Figure 2 shows, increasing the number of repetitions drastically improves the quality of the excited state energy (the difference with respect to the FCI value) improves from 25.8 mHa to 7 mHa and improves the agreement of the ground state energy with the FCI value to 2.17 mHa (from 5.6 mHa). The AC value improves from 20.2 mHa for 10 repetitions to 2.6 mHa for 30 repetitions. We extend the analysis further by increasing the number of repetitions all the way to 100, but find that the AC value remains unchanged. Thus, we instead increase the number of shots to 5000 while keeping the number of repetitions fixed at 30. Figure 3 presents our findings, with the accompanying data given in Table 1. The results indicate that as expected, one obtains better agreement with the FCI value for the AC with larger number of shots and repetitions. Furthermore, we see that the QAE result now is in much better agreement with the FCI value, differing only by 1.22 percent (5000 shots, 30 repetitions), as opposed to the 1000 shots 10 repetitions value of 13.35 percent and the 1000 shots 30 repetitions value of 3.2 percent. These percentages correspond to an energy difference ($E_{AC,FCI}-E_{AC,QAE}$) of 1.85 mHa, 20.2 mHa, and 2.6 mHa respectively. The values of AC obtained with 5000 shots and 30 repetitions, 1000 shots and 10 repetitions, and 1000 shots and 30 repetitions are 0.15354, 0.17195, and 0.15654 Ha respectively.
@@
Figure 5: Scatter plots involving fidelity and $\Delta E$ values, for 1000 shots (sub-figure (a)) and 5000 shots (sub-figure (b)) cases. The former quantity refers to $|\langle\Psi_{QAE}|\Psi_{FCI}\rangle|^{2}$, where $|\Psi_{QAE}\rangle$ is the wave function constructed out of the CI coefficients obtained using QAE and $|\Psi_{FCI}\rangle$ is the FCI wave function. On the other hand, the latter is the difference between the QAE and the FCI ground state energy results.
We turn our attention to Figure 4, where we plot the CI coefficients that we obtained from the ground state QAE calculation versus the coefficient index. We benchmark the computed CI coefficients employing QAE with those obtained from FCI. We carry out the exercise for both the 1000 and the 5000 shot cases, each with 30 repetitions. The purpose of the plot is to check the quality of the coefficients, which in turn indicates the quality of the predicted wave function. We see that the coefficients are all in reasonable agreement with their FCI counterparts. While the plot itself seems to yield almost similar set of coefficients by using 5000 shots over 1000 shots, this seemingly insignificant difference between their results is reflected in the subsequent ground state energy calculations as seen in the preceding paragraphs. Unsurprisingly, we also see the chemistry at play, where two coefficients are almost equally important (unlike the single reference scenario where only the HF state dominates). We also check the fidelity as well as the difference between AC predicted using QAE and FCI, and find them to be 99.735% and 2.6 mHa respectively for QAE calculation with 1000 shots and 30 repetitions, whereas they improve to 99.886% and 1.85 mHa respectively when we employ 5000 shots and 30 repetitions. Figures 5(a) and (b) presents scatter plots between fidelity ($\mathcal{F}$) and $\Delta E$, the difference between the QAE and the FCI ground state energy results. The fidelity is given by $|\langle\Psi_{QAE}|\Psi_{FCI}\rangle|^{2}$, where $|\Psi_{QAE}\rangle$ is the wave function constructed using the CI coefficients that QAE outputs, while $|\Psi_{FCI}\rangle$ refers to the FCI wave function. We observe from Fig. 5(a) (which has been plotted for the 1000 shots case) and (b) (the 5000 shots case) that our data unsurprisingly respects the bounds for fidelity, given by $1-\frac{E_{g,QAE}-E_{g,FCI}}{E_{e,FCI}-E_{g,FCI}}\leq\mathcal{F}\leq 1-\frac{E_{g,QAE}-E_{g,FCI}}{E_{max,FCI}-E_{g,FCI}}$, where $E_{g,FCI}$ and $E_{g,QAE}$ are the ground state energies calculated using FCI and QAE respectively, whereas $E_{e,FCI}$ and $E_{max,FCI}$ refer to the excited state of interest to our AC problem and maximum eigenvalue of $H$ obtained from FCI. For the derivation of the bound, please see Appendix A. Our results also show that most of the data points are saturating the upper bound which means a majority of the population is contained in the ground state. Furthermore, we carry out linear regression analysis on our two data sets, and find that the Pearson coefficient, $R$, is anti-correlated at $-0.84$ for the 1000 shots case and $-1.04$ in the 5000 shots case. Ideally, we would expect that as we supply more shots, the strength of anti-correlation between the two quantities, $\mathcal{F}$ and $\Delta E$, would increase, based on the bounds seen earlier. Not surprisingly, we also find that the best-case fidelities and ground state energies that we obtain using QAE for both the 1000 and the 5000 shots cases (99.735% and 99.886% respectively) obey the aforementioned inequality relation.
@@
Finally, we focus on the choice of $\lambda$ range for the problem. As mentioned in an earlier paragraph in Section II, we picked the $\lambda$ range that gave the best SAE result. For the purposes of this work, we pick a range of 0.1 Ha. However, since SAE and QAE are different approaches to search for a desired solution, we pick three such ranges of 0.1 Ha each: a range containing the HF value ($-1.77677$ Ha), which is $-1.8$ to $-1.7$ Ha, one that is immediately below and another that is immediately above it. We term the three ranges as A, B, and C respectively. We note that the range we picked for our main calculations was the one immediately above the HF value, that is, range C. Figure 7 presents our results justifying this choice. We observe that our choice of $\lambda$ range from SAE calculations were sufficiently good within a calculation involving 30 repetitions.